# On the Inherent Robustness of One-Stage Object Detection against Out-of-Distribution Data

### v1

@@ 2.1 Out-of-Distribution Detection in Object Detection
As noted in the introduction, the goal of OoD detection is to identify unknown samples in classification tasks—specifically, samples with semantics not present in the training distribution. Numerous techniques have been proposed in this area. e.g., ReAct [10], GradNorm [11], MSP [12], ODIN [13] or Energy [14]. These techniques were originally conceived for detecting OoD data in image classification tasks. In object detection tasks, however, OoD detection refers to the model’s ability to locate and identify objects in an image that do not belong to any of the classes it was trained to recognize. While traditional object detection models are trained to detect a fixed set of classes, OoD detection enables these models to recognize when an object falls outside this known set, flagging it as unknown.
@@ 2.2 Open-World Object Detection Framework
The solution proposed in [5] called ORE uses contrastive clustering, an unknown-aware proposal network and an energy-based unknown identification to address the challenges of open-world detection. This way ORE is able to predict unknown objects as “unknown” and incrementally learn new classes as their labels become available, without forgetting previously learned classes. Building on this work, [6] proposed a method called ORDER for addressing the challenges of detecting both known and unknown objects in road scenes. The authors introduce Feature-Mix, which improves unknown object detection by mixing features of known and unknown objects at the feature level. Additionally, they propose a focal regression loss to address intra-class scale variation and improve small object detection. The model is further improved through curriculum learning and is evaluated on the BDD and IDD road scene datasets.
@@
A common problem to above introduced approaches is that the module generating the proposals for the unknown objects (namely, the RPN in two-stage object detectors) is highly biased towards detecting known objects as it is the only ground truth available in the training data. To mitigate this bias, [8] substitutes the RPN with random region proposals that prevents the training from being confounded by the limited known object classes. Additionally, the authors introduce a matching score that does not penalize random proposals whose predictions do not match the known classes, encouraging the exploration of unknown objects all over the image.
Finally, the authors in [25] address exclusively the initial stage of the OWOD framework, focusing solely on the identification of unknown objects. They evaluate their method using their proposed Unknown Object Detection (UOD) benchmark, which is derived from two small subsets of the COCO dataset, augmented with annotations for unknown objects. These new annotations are generated by them and their team. The evaluation protocol involves assessing the models only on one task within these two subsets, without any incremental addition of new classes to the model. The authors also propose the so-called unknown sniffer model, which introduces a Generalized Object Confidence (GOC) score to avoid improper suppression of unknowns. Similarly, a negative energy suppression loss is used to further suppress non-object samples in the background. To obtain the best bounding box for unknown objects during inference, the paper presents a graph-based determination scheme to replace the non-maximum suppression (NMS) post-processing.
@@ 2.3 Contribution
We examine the potential of fusing OoD methods to robustly identify unknown objects. We explore the benefits of combining different OoD detection methods in the form of a detection ensemble, aiming to complement each other when deciding on the known or unknown nature of an object. Similarly to supervised learning ensembles, fusing OoD methods can yield more robust OoD detection capabilities for the overall model than every one of them in isolation.
@@ 5.2 RQ1: Which is the optimal configuration for FMap?
5.5 RQ4. How does the performance of unknown object detection algorithms implemented on single-stage models compare to that of the SOTA?
We conclude our experiments with RQ4, for which we compare the best performing FMap-based methods discovered so far to state-of-the-art methods in the UOD benchmark [25]. For the sake of readability in the reported results, we have selected the most recent and well-performing comparison baselines, and we have also included the most competitive results of the vanilla, SDR and EUL versions of FMap.
@@
| MSP | 0.470 | 0.213 | 0.314 | 0.279 | 0.359 | 0.364 | 0.055 | 0.169 | 0.190 | 0.153 | 588 | 0.135 | |||
| Mahalanobis | 0.447 | 0.129 | 0.271 | 0.309 | 0.241 | 0.351 | 0.051 | 0.149 | 0.207 | 0.116 | 604 | 0.165 | |||
| Energy score | 0.474 | 0.213 | 0.308 | 0.260 | 0.377 | 0.364 | 0.048 | 0.169 | 0.167 | 0.171 | 470 | 0.137 | |||
| ORE | 0.243 | 0.214 | 0.255 | 0.153 | 0.782 | 0.213 | 0.140 | 0.175 | 0.103 | 0.592 | 485 | 0.089 | |||
| VOS1 | 0.485 | 0.135 | 0.196 | 0.342 | 0.137 | 0.377 | 0.040 | 0.101 | 0.262 | 0.062 | 640 | 0.152 | |||
| VOS2 | 0.469 | 0.205 | 0.317 | 0.291 | 0.348 | 0.364 | 0.051 | 0.172 | 0.184 | 0.163 | 409 | 0.124 | |||
| UnSniffer | 0.464 | 0.454 | 0.479 | 0.433 | 0.535 | 0.359 | 0.150 | 0.287 | 0.222 | 0.409 | 398 | 0.175 | |||
| OW-DETR | DETR | 0.420 | 0.033 | 0.056 | 0.030 | 0.380 | 0.414 | 0.007 | 0.025 | 0.014 | 0.161 | 569 | 0.086 | ||
@@
Results are presented in Table 2. When examining the detection outcomes for known objects, it becomes evident that the mAP values for the single-stage YOLOv8 model consistently surpass those achieved by Faster R-CNN and DETR. Additionally, it is noteworthy that YOLOv8 methods exhibit generally higher precision for unknown objects. However, these methods often achieve lower recall rates, which contributes to a more balanced U-F1. Subsequently, the highest U-F1 is achieved by UnSniffer. As for the rest, FMap-based methods demonstrate enhanced robustness, as evidenced by consistently achieving lower scores in both A-OSE and WI metrics, indicative of a reduced propensity to misclassifying unknown objects as known. This robustness is attributed to our approach, as it avoids any need for specialized retraining (contrary to state-of-the-art methods, which require retraining the object detector with pseudo-labels for identifying unknown classes), significantly diminishing the likelihood of such misclassifications.

### v2

@@ 2.1 Out-of-Distribution Detection in Object Detection
As noted in the introduction, the goal of OoD detection is to identify unknown samples in classification tasks—specifically, samples with semantics not present in the training distribution. Numerous techniques have been proposed in this area. e.g., ReAct [10], GradNorm [11], MSP [12], ODIN [13], Energy [14] or Mahalanobis [15]. These techniques were originally conceived for detecting OoD data in image classification tasks. In object detection tasks, however, OoD detection refers to the model’s ability to locate and identify objects in an image that do not belong to any of the classes it was trained to recognize. While traditional object detection models are trained to detect a fixed set of classes, OoD detection enables these models to recognize when an object falls outside this known set, flagging it as unknown.
@@ 2.2 Open-World Object Detection Framework
The solution proposed in [5] called ORE uses contrastive clustering, an unknown-aware proposal network and an energy-based unknown identification to address the challenges of open-world detection. This way ORE is able to predict unknown objects as “unknown” and incrementally learn new classes as their labels become available, without forgetting previously learned classes. Building on this work, [6] proposed a method called ORDER for addressing the challenges of detecting both known and unknown objects in road scenes. The authors introduce Feature-Mix, which improves unknown object detection by mixing features of known and unknown objects at the feature level. Additionally, they propose a focal regression loss to address intra-class scale variation and improve small object detection. The model is further improved through curriculum learning and is evaluated on the BDD100K [23] and IDD [24] road scene datasets.
@@
A common problem to above introduced approaches is that the module generating the proposals for the unknown objects (namely, the RPN in two-stage object detectors) is highly biased towards detecting known objects as it is the only ground truth available in the training data. To mitigate this bias, RandBox [8] substitutes the RPN with random region proposals that prevents the training from being confounded by the limited known object classes. Additionally, the authors introduce a matching score that does not penalize random proposals whose predictions do not match the known classes, encouraging the exploration of unknown objects all over the image.
Finally, the authors in [28] address exclusively the initial stage of the OWOD framework, focusing solely on the identification of unknown objects. They evaluate their method using their proposed Unknown Object Detection (UOD) benchmark, which is derived from two small subsets of the COCO dataset, augmented with annotations for unknown objects. These new annotations are generated by them and their team. The evaluation protocol involves assessing the models only on one task within these two subsets, without any incremental addition of new classes to the model. The authors also propose the so-called unknown sniffer (UnSniffer) model, which introduces a Generalized Object Confidence (GOC) score to avoid improper suppression of unknowns. Similarly, a negative energy suppression loss is used to further suppress non-object samples in the background. To obtain the best bounding box for unknown objects during inference, the paper presents a graph-based determination scheme to replace the non-maximum suppression (NMS) post-processing.
@@ 2.3 Contribution
We examine the potential of fusing OoD methods to robustly identify unknown objects. We explore the benefits of combining different OoD detection methods in the form of a detection ensemble, aiming to complement each other when deciding on the known or unknown nature of an object. Similarly to supervised learning ensembles, fusing OoD methods can yield more robust OoD detection capabilities for the overall model than every one of them in isolation. We also show experimentally that the fusion of OoD methods with the one proposed in this work performs competitively when compared to state-of-the-art object detectors that detect unknowns by relying on specialized training procedures over pseudo-labels.
@@ 4.5 Fusion Strategies
4.6 State-of-the-art Object Detectors
In our experiments to address RQ4, we consider several avant garde methods for object detection in the presence of unknowns, including ORE [5], VOS [16], OW-DETR [25], UnSniffer [28] and RandBox [8]. For each of these comparison baselines, we use the hyperparameter configuration indicated in their publications. All these methods are within those included the UOD benchmark [28].
@@ 5.2 RQ1: Which is the optimal configuration for FMap?
5.5 RQ4. How does the performance of unknown object detection algorithms implemented on single-stage models compare to that of the state of the art?
We conclude our experiments with RQ4, for which we compare the best performing FMap-based methods discovered so far to the state-of-the-art methods listed in Subsection 4.6. For the sake of readability in the reported results, we have selected the most recent and well-performing comparison baselines, and we have also included the most competitive results of the vanilla, SDR and EUL versions of FMap.
@@
| MSP [12] | 0.470 | 0.213 | 0.314 | 0.279 | 0.359 | 0.364 | 0.055 | 0.169 | 0.190 | 0.153 | 588 | 0.135 | |||
| Mahalanobis [15] | 0.447 | 0.129 | 0.271 | 0.309 | 0.241 | 0.351 | 0.051 | 0.149 | 0.207 | 0.116 | 604 | 0.165 | |||
| Energy score [14] | 0.474 | 0.213 | 0.308 | 0.260 | 0.377 | 0.364 | 0.048 | 0.169 | 0.167 | 0.171 | 470 | 0.137 | |||
| ORE [5] | 0.243 | 0.214 | 0.255 | 0.153 | 0.782 | 0.213 | 0.140 | 0.175 | 0.103 | 0.592 | 485 | 0.089 | |||
| VOS1 [16] | 0.485 | 0.135 | 0.196 | 0.342 | 0.137 | 0.377 | 0.040 | 0.101 | 0.262 | 0.062 | 640 | 0.152 | |||
| VOS2 [16] | 0.469 | 0.205 | 0.317 | 0.291 | 0.348 | 0.364 | 0.051 | 0.172 | 0.184 | 0.163 | 409 | 0.124 | |||
| UnSniffer [28] | 0.464 | 0.454 | 0.479 | 0.433 | 0.535 | 0.359 | 0.150 | 0.287 | 0.222 | 0.409 | 398 | 0.175 | |||
| RandBox [8] | 0.790 | 0.266 | 0.329 | 0.243 | 0.507 | 0.612 | 0.052 | 0.156 | 0.112 | 0.259 | 998 | 0.109 | |||
| OW-DETR [25] | Detection Transformer | 0.420 | 0.033 | 0.056 | 0.030 | 0.380 | 0.414 | 0.007 | 0.025 | 0.014 | 0.161 | 569 | 0.086 | ||
@@
Results are presented both in Figure 13 and Table 2. Focusing on the former, we compare the above obtained non-dominated solutions for one-stage models, YOLOv8 in this case, against the state-of-the-art methods in the unknown object detection literature. It can be observed how only two of these techniques become part of the set of non-dominated solutions. UnSniffer, stands out as the method obtaining the best unknown object detection capabilities, proven by eliciting the best U-F1 score. On the other hand, RandBox achieves the best mAP of this set of solutions, while reaching acceptable U-F1. In contrast, the methodologies introduced in this study, while not achieving the highest performance across both evaluation metrics, maintain a superior equilibrium between the dual objectives. Specifically, they demonstrate strong capability for detecting unknown objects while simultaneously preserving high accuracy in identifying known entities. This is proven by the fact that three of the techniques proposed in this work are part of the non-dominated solutions in the trade-off between the performance of knowns and unknowns.
Additionally, FMap-based approaches consistently yield significantly more robust object detection models. This is evidenced by their exceptionally low A-OSE and WI metric values (see Table 2), which are the lowest among all evaluated techniques by a substantial margin—indicating a markedly reduced tendency to misclassify unknown objects as known. This robustness can be attributed to the capability of the proposed approach to eliminate the need for specialized retraining, in contrast to state-of-the-art methods that require object detectors to undergo retraining with pseudo-labels for unknown class identification. As a result, the framework facilitates the direct deployment of existing pre-trained models, thereby enhancing both practical applicability and computational efficiency.